# Optimistic Estimation of Convergence in Markov Chains with the Average-Mixing Time

### v2

@@ 2   The average-mixing time
In this section, we first recall the definition of the average-mixing time, and subsequently explore some of its basic properties. For $P\in\mathcal{W}(\mathcal{X})$, the average distance to stationarity of the chain induced from $P$ at time $t\in\mathbb{N}$ is expressed as
$d^{\sharp}(t)\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}},$
where $\pi$ is the stationary distribution of $P$. The above should be understood as the total variance distance of the chain to stationarity, averaged with respect to $\pi$ over the starting state. The limit $\lim_{t\to\infty}d^{\sharp}(t)=0$ always exists [^1] This fact follows from a connection between average-mixing time and $\beta$-mixing discussed at Section 2.3., thus for $\xi\in(0,1)$, the average-mixing time of $P$, defined (Münch and Salez, 2023) as
$\begin{split}t^{\sharp}_{\mathsf{mix}}(\xi)&\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d^{\sharp}(t)\leq\xi\right\},\end{split}$
always verifies $t^{\sharp}_{\mathsf{mix}}(\xi)<\infty$.
@@
Similar to $t_{\mathsf{mix}}$ in (2), it is tempting to introduce $t_{\mathsf{mix}}^{\sharp}\doteq t_{\mathsf{mix}}^{\sharp}(1/4)$. However, since $d^{\sharp}$ does not enjoy the same sub-multiplicativity properties as $d$ (3) (refer to (Levin et al., 2009, Lemma 4.11)), it is generally not enough to consider a fixed value of $\xi\in(1/2)$, and in this manuscript we will keep $\xi$ to be a user-fixed value.
@@ 2.1   Average-mixing versus worst-case mixing
Following their respective definitions, it holds that $d^{\sharp}(t)\leq d(t)$. Thus, for any fixed $\xi\in(0,1)$, average-mixing occurs always faster than mixing. The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2023, Remark 1), which in particular include random walks on groups [^2] A chain is called transitive when for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$.. In general however, the mixing time can be dramatically larger than the average-mixing time, even when the state space is small, as formalized below.
@@
2.2   Spectral methods under reversibility and geometric ergodicity
In this section, we endow $\mathbb{R}^{\mathcal{X}}$ with the inner product $\langle f,g\rangle_{\pi}\doteq\sum_{x\in\mathcal{X}}f(x)g(x)\pi(x)$, where $\pi$ is a positive distribution. We regard $P$ as a linear operator on the resulting Hilbert space $\ell_{2}(\pi)\doteq(\mathbb{R}^{\mathcal{X}},\langle\cdot,\cdot\rangle_{\pi})$ and further assume $P$ to be geometrically ergodic. Recall that $P\in\mathcal{W}^{\star}(\mathcal{X})$ is called $\pi$-reversible when the following detailed-balance equation holds,
which is equivalent to stating that $P$ is self-adjoint in $\ell_{2}(\pi)$. Under reversibility, the spectrum of $P$, denoted $\operatorname{Spec}(P)$, is contained in the real line. The absolute spectral gap of $P$ is then defined as
$\gamma_{\star}\doteq 1-\sup\left\{\left|\lambda\right|\colon\lambda\in\operatorname{Spec}(P),\lambda\neq 1\right\}.$
Lemma 2.1.
where $t_{\mathsf{rel}}$ is the relaxation time of $P$.
$t_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{1}{\xi\pi_{\star}},$
where $\pi_{\star}\doteq\min_{x\in\mathcal{X}}\pi(x)$, which is only informative [^4] We note in passing that the bound at Levin et al. (2009, Theorem 12.4) could be improved as $t_{\mathsf{mix}}(\xi)\leq\frac{1}{2}t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{\xi\pi_{\star}}$. when $\left|\mathcal{X}\right|<\infty$.
2.3   Connection with $\beta$-mixing
$\lim_{s\to\infty}\beta(s)=0.$
Recall that a stationary, countable-state Markov chain $X$ is $\beta$-mixing if and only if it is irreducible and aperiodic (Bradley, 2005, Theorem 3.2).
Lemma 2.2 (Davydov (1974, Proposition 1)).
$\beta(s)=\sup_{t\in\mathbb{N}}\sum_{x\in\mathcal{X}}\mu^{(t)}(x)\left\|e_{x}P^{s}-\mu^{(s+t)}\right\|_{\mathsf{TV}}.\$
$\begin{split}\beta(s)=\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{s}-\pi\right\|_{\mathsf{TV}}=d^{\sharp}(s),\\ \end{split}$
thus the average-mixing time is essentially the “stationary $\beta$-mixing time” of the chain555 Similarly, the worst-case mixing time can be connected to the notion of $\phi$-mixing (Davydov, 1974). . Note that with the convention $P^{0}=I$, we obtain
$\beta(0)=1-\left\|\pi\right\|_{2}^{2},$
which depends only on the stationary properties of the chain. When there exists $\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{1}/s^{b}$, we say that the process $\beta$-mixes algebraically or polynomially. When there exists $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially when $b\neq 1$, and exponentially when $b=1$. In Section 6, we will introduce a class of Markov transition operators, whose flavor of mixing (algebraic, sub-exponential, exponential) can be fully controlled. Notably (refer to the proof of Theorem 3.2), when $P$ is uniformly ergodic with mixing time $t_{\mathsf{mix}}$, it holds that
$\beta(s)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$ (4)
hence the process is exponentially $\beta$-mixing with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$. However, nothing a priori precludes the $\beta$-mixing rate from being superior to (4). The connection between the average-mixing time and $\beta$-mixing provides additional motivation for the problem of estimating the $\beta$-mixing coefficients, whose knowledge is known to lead to powerful decoupling techniques (Yu, 1994). The problem of inferring mixing coefficients from a single trajectory of observations has recently received attention from the community. In the most recent work (Khaleghi and Lugosi, 2023), consistent estimators for $\beta$- and $\alpha$-mixing coefficients were proposed. In Section 3, we will provide more explicit and finer finite-sample bounds for the estimation problem in the specialized Markovian setting. A regime of particular interest to us is when the chain $\beta$-mixes much faster than $\Theta(t_{\mathsf{mix}})$. Henceforth in this manuscript, when they coincide under stationarity, we may use the notations $\beta$ and $d^{\sharp}$ interchangeably.
2.4   Bounding the deviation of additive functionals evaluated on Markov chains
@@
We now illustrate how the average-mixing time, through its connection to $\beta$-mixing naturally appears as a quantity of interest when bounding deviations of functions evaluated on a trajectory sampled from a Markov chain. More formally, let $X_{1},\dots,X_{n}$ be a Markov chain over a countable space $\mathcal{X}$, and whose dynamics are governed by $P\in\mathcal{W}(\mathcal{X})$ with stationary distribution $\pi$. For simplicity, we consider a real bounded function $f$ on $\mathcal{X}$, which is centered in the sense where its stationary expected value vanishes $\mathbb{E}_{\pi}f=0$. The ergodic theorem states that the empirical mean taken on an infinite trajectory will similarly vanish,
@@
In practice however, we are more likely to be interested in controlling the deviation of the sample mean from its expected value in a non-asymptotic setting.
@@
Let $f\colon\mathcal{X}\to[-1,1]$ with $\mathbb{E}_{\pi}f=0$, let $\varepsilon,\delta\in(0,1)$, and let $t^{\sharp}_{\mathsf{mix}}\colon(0,1)\to\mathbb{N}$. For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, and average-mixing time bounded from above by $t^{\sharp}_{\mathsf{mix}}$, the following statements hold.
@@
    with
    $\xi(\varepsilon,\delta)\doteq\frac{\delta\varepsilon^{2}}{16\log\left(\frac{4}{\delta}\right)},$
    it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$. The above upper on the path length can be further streamlined as
@@
We adopt the traditional method as described in Yu (1994). This approach combines a blocking technique, credited to Bernstein (1927) in the context of $m$-dependent processes, together with an application of Hoeffding’s inequality to control the deviation probability of a sum of independent observations. We further optimize for the block size and reformulate the results to obtain an expression involving the average-mixing time under various rates of $\beta$-mixing. Refer to Section 5.1 for the proof. ∎
@@
As a consequence of Lemma 2.3, in both the sub-exponential and polynomial settings, provided an assumption on the convergence order $b$, estimates on the average-mixing time to proximity $\Theta(\delta)$ yield corresponding deviation bounds. In the uniformly ergodic setting, corresponding to the assumption of the exponential setting ($b=1$), we obtain a deviation inequality involving the average-mixing time in lieu of the mixing time (Chung et al., 2012, Theorem 3.1). The lemma generalizes to non-centered functions, lower deviations and arbitrarily bounded functions. It is also noteworthy that in the sub-exponential setting, the convergence order $b$ solely affects the constant in the sample complexity. Our formulation above serves as additional motivation for analyzing and inferring the average-mixing time.
3   Estimation from a single trajectory
In this section, we analyze the estimation of $\beta$-mixing coefficients and the average-mixing time from a single trajectory of observations.
3.1   Estimation of $\beta$-mixing coefficients under general ergodicity condition
In this section we analyse plug-in estimators for $\beta$-mixing coefficients. We recall that in this general setting, the mixing time can be infinite, and the $\beta$-mixing coefficients can decay at arbitrary slow rate. We will focus on exponential, sub-exponential and polynomial rates of convergence. Our analysis relies on stationary skipped trajectories, where for $s\in\mathbb{N},s<n$, we denote

### v4

@@ 2  The average-mixing time
In this section, we first recall the definition of the average-mixing time and then explore some of its basic properties. For a $\pi$-stationary transition matrix $P\in\mathcal{W}(\mathcal{X})$ and for a proximity parameter $\xi\in(0,1)$, the average-mixing time of $P$ is defined (Münch and Salez, 2023) as
$\begin{split}t^{\sharp}_{\mathsf{mix}}(\xi)&\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{\beta(t)\leq\xi\right\},\end{split}$
where $\beta\colon\mathbb{N}\to[0,1]$ is a sequence defined by
$\beta(t)\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}}.$ (4)
Observe that instead of considering the distance to stationarity from the worst-case starting state, as the distance $d$ does, $\beta$ takes an average distance over starting states with respect to the long term probability of being in each state. This choice naturally reduces the contribution of states that are difficult to reach, resulting in a more optimistic convergence criterion.
@@
Similar to the definition of $t_{\mathsf{mix}}$ in (2), one may wish to introduce $t_{\mathsf{mix}}^{\sharp}\doteq t_{\mathsf{mix}}^{\sharp}(1/4)$. However, since $\beta$ does not enjoy the same sub-multiplicativity properties as $d$ —refer to (3) and (Levin et al., 2009, Lemma 4.11)—it is generally not enough to consider an arbitrary value of $\xi\in(1/2)$, and in this manuscript we will keep $\xi$ as a user-fixed value.
Remark 2.2 (Average-mixing time is stationary $\beta$-mixing time).
The astute reader will have noticed that we choose $\beta$ to denote the average distance to stationarity instead of $d^{\sharp}$, as in (Münch and Salez, 2023). This choice is motivated by our observation that the average-mixing time corresponds to the stationary $\beta$-mixing time of the Markov chain, which we discuss in detail in Section B. The sequence $\beta$ defined at (4) will henceforth be referred to as the sequence of (stationary) $\beta$-mixing coefficients.
An immediate consequence of Remark 2.2 is that the limit $\lim_{t\to\infty}\beta(t)=0$ always exists (Bradley, 2005), or in other words, $t^{\sharp}_{\mathsf{mix}}(\xi)<\infty$ for any $\xi\in(0,1)$ and any ergodic $P\in\mathcal{W}(\mathcal{X})$ over countable $\mathcal{X}$. However, the rate of convergence of the sequence of $\beta$-mixing coefficients can be arbitrarily slow.
Definition 2.1.
Rates of convergence for $\beta$-mixing.
  1. 1.
    (Sub-exponential, exponential) When there exists $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$ and $b\in(0,1]$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially when $b\neq 1$, and exponentially when $b=1$.
  2. 2.
    (Polynomial) When there exists $\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{1}/s^{b}$, we say that the process $\beta$-mixes algebraically or polynomially.
Notably, when $P$ is uniformly ergodic with mixing time $t_{\mathsf{mix}}$, it holds that
$\beta(s)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$ (5)
hence the process is exponentially $\beta$-mixing with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$. However, nothing a priori precludes the $\beta$-mixing rate from being much faster than (5).
@@ 2.1  Average-mixing versus worst-case mixing
From their respective definitions, it is immediate that $\beta(t)\leq d(t)$. In other words, for any fixed $\xi\in(0,1)$, average-mixing never occurs slower than mixing. The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2023, Remark 1), which are chains such that for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$ (Levin et al., 2009, Section 2.6.2). Transitive chains include, but are not limited to, random walks on groups (McKay and Praeger, 1996). For non-transitive chains, the mixing time can be dramatically larger than the average-mixing time, even when the state space is small, as made precise below.
@@
In fact, as we will see in the next section, a regime of particular interest to us is when the chain $\beta$-mixes much faster than $\Theta(t_{\mathsf{mix}})$. We begin by recalling the well-studied class of birth and death Markov chains, which will serve as our running example.
Example 2.1 (Birth and death Markov chains).
We let $\mathcal{X}=\mathbb{N}$, and for three sequences $u,v,w\in[0,1]^{\mathbb{N}}$ such that $u(1)=0$ and $\forall n\in\mathbb{N},u(n)+v(n)+w(n)=1$, we define the transition operator,
$P_{u,v,w}=\begin{pmatrix}v(1)&w(1)&0&0&0&\ldots\\ u(2)&v(2)&w(2)&0&0&\vdots\\ 0&u(3)&v(3)&w(3)&0&\ddots\\ 0&0&u(4)&v(4)&\ddots&\ddots\\ \vdots&\ldots&\ddots&\ddots&\dots&\ddots\\ \end{pmatrix}$.
Due to their ubiquity, ergodicity properties of birth and death chains have been extensively studied in the literature (Miclo, 1999; Ding et al., 2010; Kovchegov, 2010; Chen and Saloff-Coste, 2013). In the special case where $v(1)=0$ and for $n\geq 2$ the sequences $u,v,w$ are constant with $u>w$, it can be shown that the family of chains
$P_{u,v,w}=\begin{pmatrix}0&1&0&0&0&\ldots\\ u&v&w&0&0&\vdots\\ 0&u&v&w&0&\ddots\\ 0&0&u&v&\ddots&\ddots\\ \vdots&\ldots&\ddots&\ddots&\dots&\ddots\\ \end{pmatrix}$
is geometrically ergodic (Meitz and Saikkonen, 2021; Kovchegov, 2010), that is, there exists $M\colon\mathcal{X}\to\mathbb{R}_{+}$ and $\rho\in(0,1)$ such that for any $t\in\mathbb{N}$ and any $n_{0}\in\mathcal{X}$,
$\left\|e_{n_{0}}P_{u,v,w}^{t}-\pi\right\|_{\mathsf{TV}}\leq M(n_{0})\rho^{t}.$ (6)
In particular, the geometric rate of the above family is given by
$\rho=\max\left\{v+2\sqrt{uw},\frac{w}{w+v}\right\}.$
While geometric ergodicity is weaker than uniform ergodicity in as much as the distance to the stationary distribution depends on the starting state, convergence remains exponential and leads to a powerful convergence analysis based on spectral methods (refer also to Section 2.3.1). But we now turn our attention to the even less favorable case of sub-geometric convergence, where the aforementioned spectral methods break down. This more challenging setting can arise even in for the seemingly simple class of birth and death chains in this example. Indeed, let us borrow the following instantiation of a Chebyshev–type random walk, proposed by Kovchegov and Michalowski (2013, Section 4). Let $\theta>0$ and $\lambda\geq\frac{2\theta^{2}}{(1+\theta)(1+3\theta)}$. We define
[…]
while for $n\geq 2$,
[…]
and v(n)=1u(n)w(n)v(n)=1-u(n)-w(n). For convenience, we denote $P_{\theta,\lambda}=P_{u,v,w}$ for the above-defined sequences. It can be verified that the stationary distribution of $P_{\theta,\lambda}$ satisfies for any $n\in\mathbb{N}$,
$\pi_{\theta}(n)=\frac{2(1+\theta)\theta}{(1+(2n-1)\theta)(1+(2n-3)\theta)}\asymp\frac{1}{n^{2}}.$
One interesting feature is that the chain does not have a spectral gap—or in other words, it “[lies] outside the scope of geometric ergodicity theory” (Kovchegov and Michalowski, 2013). In fact, for the random walk originating at state $n_{0}\in\mathbb{N}$, it holds (Kovchegov and Michalowski, 2013, Theorem 2) that for any $t\in\mathbb{N}$,
$\frac{c}{\sqrt{t}}\leq\left\|e_{n_{0}}P_{\theta,\lambda}^{t}-\pi_{\theta}\right\|_{\mathsf{TV}}\leq\frac{Cn_{0}\log(t+n_{0}+2)}{\sqrt{t}},$
where $c$ is a constant depending on $\theta$, $\lambda$ and the starting state $n_{0}$, while $C$ is a constant which depends only on $\theta$ and $\lambda$. Let $\xi\in(0,1)$ and denote
$n_{\star}=n_{\star}(\theta,\xi/2)=\min_{n\in\mathbb{N}}\left\{\sum_{k=n+1}^{\infty}\pi_{\theta}(k)<\xi/2\right\}.$
We can thus bound the $\beta$-mixing coefficient as follows,
[…]
where $C^{\prime}$ is a constant which depends only on $\theta$ and $\lambda$ and for any $k\in\mathbb{N}$, $H_{k}$ denotes the $k$th harmonic number which is bounded from above as $H_{k}\leq 1+\log(k)$. Furthermore, one can verify that $n_{\star}\leq 2+\sqrt{(1+\theta)/(\xi\theta)}$. As a result, there exists a constant $C^{\prime\prime}$ which depends only on $\theta$ and $\lambda$ such that,
[…]
and we obtain that
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\frac{1}{\xi^{2}}\operatorname{polylog}(1/\xi),$
where $\operatorname{polylog}$ is a poly-logarithmic function of its argument, which may also depend on $\theta$ and $\lambda$. Our example demonstrates how the contributions to $\beta(s)$ starting from the tail of the stationary distribution tend to be largely absorbed.
2.2  Deviation of additive functionals of Markov chains
@@
We now illustrate how the average-mixing time, through its connection to $\beta$-mixing naturally appears as a quantity of interest when bounding deviations of functions evaluated on a trajectory sampled from a Markov chain. More formally, let $X_{1},\dots,X_{n}$ be a Markov chain over a countable space $\mathcal{X}$, and whose dynamics are governed by $P\in\mathcal{W}(\mathcal{X})$ with stationary distribution $\pi$. For simplicity, we consider a real bounded function $f$ on $\mathcal{X}$, which is centered in the sense where its stationary expected value vanishes, that is $\mathbb{E}_{\pi}f=0$. The ergodic theorem states that the empirical mean taken on an infinite trajectory will similarly vanish,
@@
In practice however, we are more likely to be interested in controlling the deviation of the sample mean from its expected value non-asymptotically. The following lemma illustrates how to recover a finite sample bound on the deviation probability, given a bound on the average-mixing time and the rate of convergence of the sequence of $\beta$-mixing coefficients.
@@
Let $f\colon\mathcal{X}\to[-1,1]$ with $\mathbb{E}_{\pi}f=0$, let $\varepsilon,\delta\in(0,1)$. For any $\pi$-stationary ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, and average-mixing time $t^{\sharp}_{\mathsf{mix}}\colon(0,1)\to\mathbb{N}$, the following statements hold.
@@
    with ξ(ε, δ) δε216 log(4δ), it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$. The above upper on the path length can be further streamlined as
@@
We adopt a traditional method described in Yu (1994). This approach combines a blocking technique, credited to Bernstein (1927) in the context of $m$-dependent processes, together with an application of Hoeffding’s inequality to control the deviation probability of a sum of independent observations. We further optimize for the block size and reformulate the results to obtain an expression involving the average-mixing time under various rates of $\beta$-mixing. Refer to the supplementary material for the complete proof. ∎
@@
As a consequence of Lemma 2.1, in both the sub-exponential and polynomial settings, provided an assumption on the convergence order $b$, estimates on the average-mixing time to proximity $\Theta(\delta)$ yield corresponding deviation bounds. In the uniformly ergodic setting, corresponding to the assumption of the exponential setting ($b=1$), we obtain a deviation inequality involving the average-mixing time in lieu of the mixing time (Chung et al., 2012, Theorem 3.1), (Paulin, 2015, Corollary 2.10). The lemma generalizes to non-centered functions, lower deviations and arbitrarily bounded functions. It is also noteworthy that in the sub-exponential setting, the convergence order $b$ solely affects the constant in the sample complexity. Our formulation above serves as motivation for analyzing and inferring the average-mixing time.
Remark 2.3.
The astute reader will observe that, in the uniformly ergodic setting, the logarithmic dependence in the confidence parameter $\delta$ is slightly sharper for bounds involving the worst-case mixing time. (Chung et al., 2012, Theorem 3.1), (Paulin, 2015, Corollary 2.10). This is a consequence of our proof technique relying on $\beta$-mixing.
2.3  Under geometric ergodicity
In this section, we briefly examine the average-mixing time under geometric ergodicity, first using spectral methods under time-reversibility, and then through more broadly applicable Lyapunov-type methods.
2.3.1 Spectral methods under reversibility
Spectral methods are known to provide a direct and concise framework for analyzing the worst-case mixing time. In this section, we briefly examine their applicability to the study of the average-mixing time. We endow $\mathbb{R}^{\mathcal{X}}$ with the inner product $\langle f,g\rangle_{\pi}\doteq\sum_{x\in\mathcal{X}}f(x)g(x)\pi(x)$, where $\pi$ is a positive distribution. We regard $P$ as a linear operator on the resulting Hilbert space $\ell_{2}(\pi)\doteq(\mathbb{R}^{\mathcal{X}},\langle\cdot,\cdot\rangle_{\pi})$ and further assume $P$ to be geometrically ergodic—refer to (6). Recall that $P\in\mathcal{W}^{\star}(\mathcal{X})$ is called $\pi$-reversible when the following detailed-balance equation holds,
which is equivalent to stating that $P$ is self-adjoint in $\ell_{2}(\pi)$. Under reversibility, the spectrum of $P$, denoted $\operatorname{Spec}(P)$, is contained in the real line. The absolute spectral gap of $P$ is defined as
$\gamma_{\star}\doteq 1-\sup\left\{\left|\lambda\right|\colon\lambda\in\operatorname{Spec}(P),\lambda\neq 1\right\}.$ (7)
Lemma 2.2.
where $t_{\mathsf{rel}}$ is the relaxation time of $P$ defined by $t_{\mathsf{rel}}=1/\gamma_{\star}$ with $\gamma_{\star}$ the absolute spectral gap of $P$ defined in (7).
$t_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{1}{\xi\pi_{\star}},$
where $\pi_{\star}\doteq\min_{x\in\mathcal{X}}\pi(x)$, which is only informative when $\left|\mathcal{X}\right|<\infty$.
2.3.2 Lyapunov methods
Following a classical Banach space structure, the definitions of the $V$-norm of a function $f$ on $\mathcal{X}$ and a signed measure $\mu$ on $\mathcal{X}$ are given (Douc et al., 2018, Definition D.3.1) respectively as ,
$\left\|f\right\|_{V}\doteq\sup_{x\in\mathcal{X}}\frac{\left|f(x)\right|}{V(x)}\qquad\left\|\mu\right\|_{V}\doteq\sum_{x\in\mathcal{X}}\left|\mu(x)\right|V(x)=\sup_{\left\|f\right\|_{V}\leq 1}\left|\sum_{x\in\mathcal{X}}f(x)\mu(x)\right|.$
We say that $P$ satisfies $V$-geometrical ergodicity (Hervé and Ledoux, 2020) when there exists and $\rho_{V}\in(0,1)$ such that for all $x\in\mathcal{X}$ and all $t\in\mathbb{N}$,
$\left\|e_{x}P^{t}-\pi\right\|_{V}\leq V(x)\rho_{V}^{t}.$
In particular, $P$ is known to be $V$-geometrically ergodic under irreducibility, aperiodicity, drift and minorization conditions (Hervé and Ledoux, 2020). In this case, observe that we can bound the stationary $\beta$-coefficient as follows,
[…]
As a result,
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\cfrac{\log\left\|\pi\right\|_{V}+\log\frac{1}{2\xi}}{\log\frac{1}{\rho_{V}}},$
which is finite whenever $\left\|\pi\right\|_{V}=\mathbb{E}_{\pi}\left[V\right]$ is finite.
3  Estimation of average convergence from a single trajectory
To apply the results in Section 2, one needs to have an upper bound on the average-mixing time. However, such bound may not be known a priori, and may need to be estimated from the data. In this section, we consider the problem of estimating the average-mixing time $t^{\sharp}_{\mathsf{mix}}$ from a single trajectory of observations. Specifically, for a user-fixed proximity parameter $\xi\in(0,1)$ and a confidence parameter $\delta\in(0,1)$, our goal is to construct an interval $I_{\delta}\subset[1,\infty)$, such that given a trajectory $X_{1},\dots,X_{n}$ sampled from an unknown transition matrix $P$, and for $n$ sufficiently large, it holds that $t_{\mathsf{mix}}^{\sharp}(\xi)\in I_{\delta}$ with probability at least $1-\delta$.
We first analyze the intermediary problem of estimating the $\beta$-mixing coefficients of the process. We subsequently show how to convert such results into estimation procedures for the average-mixing time both in the uniformly ergodic and non-uniformly ergodic settings (refer to Figure 1).
[…]
Figure 1: Logical flow between the estimation results of Section 3. MAD: Mean Absolute Deviation; PAC: Probably Approximately Correct.
3.1  Estimation of $\beta$-mixing coefficients
We begin by analyzing a sequence of plug-in estimators for the $\beta$-mixing coefficients. Our analysis relies on stationary skipped trajectories, where for $s\in\mathbb{N},s<n$, we denote